# Exploring Homogeneous and Heterogeneous Consistent Label Associations for Unsupervised Visible-Infrared Person ReID

### v2

@@ 3.2 Modality-Invariant Representation Learning
Method Venue SYSU-MM01 RegDB All-search Indoor-search Visible-to-Infrared Visible-to-Infrared R1 mAP mINP R1 mAP mINP R1 mAP mINP R1 mAP mINP Supervised VI-ReID methods AGW [51] TPAMI-21 47.50 47.65 35.30 54.17 62.97 59.23 70.05 66.37 50.19 70.49 65.90 51.24 VCD+VML [40] CVPR-21 60.02 58.80 - 66.05 72.98 - 73.2 71.6 - 71.8 70.1 - MPANet [45] CVPR-21 70.58 68.24 - 76.74 80.95 - 82.8 80.7 - 83.7 80.9 - MAUM [28] CVPR-22 71.68 68.79 - 76.97 81.94 - 87.87 85.09 - 86.95 84.34 - CIFT [23] ECCV-22 74.08 74.79 - 81.82 85.61 - 91.96 92.00 - 90.30 90.78 - SEFEL [20] CVPR-23 77.12 72.33 - 82.07 82.95 - 91.07 85.23 - 92.18 86.59 - PartMix [20] CVPR-23 77.78 74.62 - 81.52 84.38 - 84.93 82.52 - 85.66 82.27 - Unsupervised single-modality ReID methods MMT [14] ICLR-20 21.47 21.53 11.50 22.79 31.50 27.66 25.68 26.51 19.56 24.42 25.59 18.66 Cluster-Contrast [11] arXiv-21 20.16 22.00 12.97 23.33 34.01 30.88 11.76 13.88 9.94 11.14 12.99 8.99 ICE [3] ICCV-21 20.54 20.39 10.24 29.81 38.35 34.32 12.98 15.64 11.91 12.18 14.82 10.6 PPLR [8] CVPR-22 11.98 12.25 4.97 12.71 20.81 17.61 10.30 11.94 8.10 10.39 11.23 7.04 Unsupervised VI-ReID methods H2H [24] TIP-21 25.49 25.16 - - - - 13.91 12.72 - 14.11 2.29 - H2H(AGW) [24] TIP-21 30.15 29.40 - - - - 23.81 18.87 - - - - H2H(AGW) w/ CMRR [24] TIP-21 45.47 47.99 - - - - 35.18 36.46 - - - - OTLA [42] ECCV-22 29.9 27.1 - 29.8 38.8 - 32.9 29.7 - 32.1 28.6 - ADCA [47] MM-22 45.51 42.73 28.29 50.60 59.11 55.17 67.20 64.05 52.67 68.48 63.81 49.62 ADCA(AGW) [47] MM-22 50.90 45.70 29.12 51.39 59.82 56.08 66.62 63.47 - 67.29 62.98 - CHCR (AGW) [34] T-CSVT-23 47.72 45.34 - - - - 68.18 63.75 - 69.96 65.87 - DOTLA (AGW) [6] MM-23 50.36 47.36 32.40 53.47 61.73 57.35 85.63 76.71 61.58 82.91 74.97 58.60 MBCCM [5] MM-23 53.14 48.16 32.41 55.21 61.98 57.23 83.79 77.87 65.04 82.82 76.74 61.73 CCLNet (CLIP) [4] MM-23 54.03 50.19 - 56.68 65.12 - 69.94 65.53 - 70.17 $6$ PGM(AGW) [46] CVPR-23 57.27 51.78 34.96 56.23 62.74 58.13 69.48 65.41 - 69.85 65.17 - ${}^{\ast}$ [48] ICCV-23 60.95 56.99 41.85 64.22 69.49 64.81 73.91 70.23 58.88 75.00 69.94 56.21 Ours - 65.03 58.62 42.77 65.35 71.24 66.60 91.50 83.73 69.13 89.08 80.88 64.03 Ours w/ CMRR [24] Ours + TIP-21 71.39 67.68 55.48 69.75 74.53 70.16 92.77 93.86 93.04 90.34 91.91 91.18
@@ 3.3 Online Cross-memory Label Refinement
To overcome this issue, we utilize the predictions from the intra-modality memory banks to refine the predictions from the cross-modality memory bank. Several reasons support the adoption of such refinement strategies: (1) The intra-modality memory banks are updated by features within their respective modalities, this update occurs at a slower rate compared to the cross-modality memory bank, resulting in more stable predictions. (2) The intra-modality and cross-modality memory banks are expected to output consistent predictions. The OCLR loss for visible modality can be formulated as follows:
$\mathcal{L}^{v}_{OCLR}$ $=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\mathbf{M}^{a},\tau),P(\mathbf{f}_{i}^{v}|\tilde{\mathbf{M}}^{v},\frac{1}{5}\tau))$ (19)
$+\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\mathbf{M}^{a},\tau),\mathbf{P}(\mathbf{f}_{i}^{v}|\hat{\mathbf{M}}^{r},\frac{1}{5}\tau)).$
As in [1, 32], we sharpen the target prediction more than the source’s to encourage entropy minimization. The setting of the sharpen ratio follows ProtoCon [32]. In the visible modality, our OCLR operated between the visible stream and the augmented stream [47], enhancing constraints on self-consistency. The OCLR loss for infrared modality $\mathcal{L}^{r}_{OCLR}$ can be formulated similarly.
3.4 Optimization
The total training loss can be formulated as follows:
@@ 3.4 Optimization
where $\mathcal{L}_{IM}$, $\mathcal{L}_{CM}$, $\mathcal{L}^{v}_{OCLR}$ and $\mathcal{L}^{r}_{OCLR}$ are described in detail in Sec.3.2 and Sec.3.3. The training process is summarized in the supplementary meterial.
@@ 4 Experiments
Index Method SYSU-MM01 RegDB All-search Indoor-search Visible-to-Infrared Visible-to-Infrared R1 mAP mINP R1 mAP mINP R1 mAP mINP R1 mAP mINP 1 Baseline 41.76 38.57 24.32 45.83 54.44 50.00 50.58 48.78 35.62 51.17 47.24 33.80 The following methods are all based on the MULT module 2 Baseline + $\mathcal{L}_{CM}$ 60.43 53.98 37.28 58.79 64.61 59.29 76.31 67.97 50.62 78.11 66.55 45.64 3 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{CM}$ 61.48 54.43 38.02 60.73 66.32 61.52 76.21 67.94 51.40 74.95 65.34 45.46 4 Baseline + $\mathcal{L}_{OCLR}$ 51.88 47.37 31.39 52.63 59.44 55.29 86.80 80.36 66.44 86.84 79.29 61.36 5 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{OCLR}$ 61.43 56.08 40.59 64.67 70.43 66.46 88.59 82.18 68.84 88.01 80.57 64.22 6 Baseline + $\mathcal{L}_{CM}$ + $\mathcal{L}_{OCLR}$ 61.16 56.21 40.92 62.18 68.31 63.94 88.64 82.05 68.85 87.62 80.82 65.26 7 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{CM}$ + $\mathcal{L}_{OCLR}$ 65.03 58.62 42.77 65.35 71.24 66.60 91.50 83.73 69.13 89.08 80.88 64.03

### v4

@@ 3.2 Alternative Modality-Invariant Representation Learning
Method Venue SYSU-MM01 (Single-shot) RegDB All-search Indoor-search Visible-to-Infrared Visible-to-Infrared R1 mAP mINP R1 mAP mINP R1 mAP mINP R1 mAP mINP Supervised VI-ReID methods Zero-PadWu et al (2017) ICCV-17 14.80 15.95 - 20.58 26.92 - 17.75 18.90 - 16.63 17.82 - AlignGANMao et al (2017) ICCV-19 42.4 40.7 - 45.9 54.3 - 57.9 53.6 - 56.3 53.4 - cm-SSFT Lu et al (2020) CVPR-20 47.7 54.1 - - - - 72.3 72.9 - 71.0 71.7 - DDAG Ye et al (2020) ECCV-21 54.75 53.02 39.62 61.02 67.98 62.61 69.34 63.46 49.24 68.06 61.80 48.62 AGW Ye et al (2021b) TPAMI-21 47.50 47.65 35.30 54.17 62.97 59.23 70.05 66.37 50.19 70.49 65.90 51.24 VCD+VML Tian et al (2021) CVPR-21 60.02 58.80 - 66.05 72.98 - 73.2 71.6 - 71.8 70.1 - CA Ye et al (2021a) ICCV-21 69.88 66.89 53.61 76.26 80.37 76.79 85.03 79.14 65.33 84.75 77.82 61.56 MPANet Wu et al (2021) CVPR-21 70.58 68.24 - 76.74 80.95 - 82.8 80.7 - 83.7 80.9 - MAUM Liu et al (2022) CVPR-22 71.68 68.79 - 76.97 81.94 - 87.87 85.09 - 86.95 84.34 - CIFT Li et al (2022b) ECCV-22 74.08 74.79 - 81.82 85.61 - 91.96 92.00 - 90.30 90.78 - DEENZhang and Wang (2023) CVPR-23 74.7 71.8 - 80.3 83.3 - 91.1 85.1 - 89.5 83.4 - SEFEL Feng et al (2023) CVPR-23 77.12 72.33 - 82.07 82.95 - 91.07 85.23 - 92.18 86.59 - PartMix Kim et al (2023) CVPR-23 77.78 74.62 - 81.52 84.38 - 84.93 82.52 - 85.66 82.27 - Unsupervised single-modality ReID methods SPCL Ge et al (2020b) NIPS-20 18.37 19.39 10.99 26.83 36.42 33.05 13.59 14.86 10.36 11.70 13.56 10.09 MMT Ge et al (2020a) ICLR-20 21.47 21.53 11.50 22.79 31.50 27.66 25.68 26.51 19.56 24.42 25.59 18.66 Cluster-Contrast Dai et al (2023) arXiv-21 20.16 22.00 12.97 23.33 34.01 30.88 11.76 13.88 9.94 11.14 12.99 8.99 ICE Chen et al (2021) ICCV-21 20.54 20.39 10.24 29.81 38.35 34.32 12.98 15.64 11.91 12.18 14.82 10.6 PPLR Cho et al (2022) CVPR-22 11.98 12.25 4.97 12.71 20.81 17.61 10.30 11.94 8.10 10.39 11.23 7.04 ISE Zhang et al (2022b) CVPR-22 20.21 18.93 8.54 14.22 24.62 21.74 16.12 16.99 13.24 10.83 13.66 10.71 Unsupervised VI-ReID methods H2H Liang et al (2021) TIP-21 25.49 25.16 - - - - 13.91 12.72 - 14.11 12.29 - H2H(AGW) Liang et al (2021) TIP-21 30.15 29.40 - - - - 23.81 18.87 - - - - H2H(AGW) w/ CMRR Liang et al (2021) TIP-21 45.47 47.99 - - - - 35.18 36.46 - - - - OTLA Wang et al (2022) ECCV-22 29.9 27.1 - 29.8 38.8 - 32.9 29.7 - 32.1 28.6 - ADCA Yang et al (2022) MM-22 45.51 42.73 28.29 50.60 59.11 55.17 67.20 64.05 52.67 68.48 63.81 49.62 ADCA(AGW) Yang et al (2022) MM-22 50.90 45.70 29.12 51.39 59.82 56.08 66.62 63.47 - 67.29 62.98 - CHCR (AGW) Pang et al (2023) T-CSVT-23 47.72 45.34 - - - - 68.18 63.75 - 69.96 65.87 - DOTLA (AGW) Cheng et al (2023b) MM-23 50.36 47.36 32.40 53.47 61.73 57.35 85.63 76.71 61.58 82.91 74.97 58.60 MBCCM Cheng et al (2023a) MM-23 53.14 48.16 32.41 55.21 61.98 57.23 83.79 77.87 65.04 82.82 76.74 61.73 MIMR Pang et al (2024) KBS-24 46.56 45.88 - 52.26 60.93 - 68.76 64.33 - 68.76 63.83 - CCLNet (CLIP) Chen et al (2023) MM-23 54.03 50.19 - 56.68 65.12 - 69.94 65.53 - 70.17 $6$ PGM(AGW) Wu and Ye (2023) CVPR-23 57.27 51.78 34.96 56.23 62.74 58.13 69.48 65.41 - 69.85 65.17 - GUR Yang et al (2023) ICCV-23 60.95 56.99 41.85 64.22 69.49 64.81 73.91 70.23 58.88 75.00 69.94 56.21 Ours - 64.77 59.23 43.46 65.34 71.46 67.83 89.95 82.09 67.29 90.78 82.25 65.38 Ours w/ CMRR Liang et al (2021) Ours + TIP-21 71.23 67.54 54.48 69.06 75.06 71.36 91.50 92.51 92.03 90.53 91.49 91.27
@@ 3.3 Online Cross-memory Label Refinement
To overcome this issue, we utilize the predictions from the intra-modality memory banks to refine the predictions from the cross-modality memory bank. Several reasons support the adoption of such refinement strategies: (1) The intra-modality memory banks are updated by features within their respective modalities, this update occurs at a slower rate compared to the cross-modality memory bank, resulting in more stable predictions. (2) The intra-modality and cross-modality memory banks are expected to output consistent predictions. In the visible modality, we adopt mutual refinement between the extracted features $\mathbf{f}_{i}^{v}$ from visible images and $\mathbf{f}_{i}^{a}$ from images with CA augmentation Ye et al (2021a) to enhance constraints on their self-consistency. CA is a widely-used data augmentation strategy in VI-ReID Yang et al (2022); Cheng et al (2023a); Wu and Ye (2023), which randomly selects one channel in a visible image and expands it into a three-channel gray image. The augmented images build an intermediate modality between visible and infrared to enhance training. The OCLR loss for visible modality in V-based training mode can be formulated as follows:
$\mathcal{L}^{v}_{OCLR}$ $=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{v}|\mathbf{M}^{c},\tau),\mathbf{P}(\mathbf{f}_{i}^{a}|\tilde{\mathbf{M}}^{v},\frac{1}{5}\tau))$ (27)
$+\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{a}|\mathbf{M}^{c},\tau),\mathbf{P}(\mathbf{f}_{i}^{v}|\hat{\mathbf{M}}^{r},\frac{1}{5}\tau)).$
As in Caron et al (2021); Nassar et al (2023), we sharpen the target prediction more than the source’s to encourage entropy minimization. The setting of the sharpen ratio follows ProtoCon Nassar et al (2023). The OCLR loss for infrared modality $\mathcal{L}^{r}_{OCLR}$ in V-based mode can be formulated similarly:
$\mathcal{L}^{r}_{OCLR}$ $=\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{r}|\mathbf{M}^{c},\tau),\mathbf{P}(\mathbf{f}_{i}^{r}|\tilde{\mathbf{M}}^{v},\frac{1}{5}\tau))$ (28)
$+\frac{1}{B}\sum^{B}_{i=1}\mathcal{L}_{ce}(\mathbf{P}(\mathbf{f}_{i}^{r}|\mathbf{M}^{c},\tau),\mathbf{P}(\mathbf{f}_{i}^{r}|\hat{\mathbf{M}}^{r},\frac{1}{5}\tau)).$
3.4 The Overall Objective Function
The overall training loss can be formulated as follows:
@@ 3.4 The Overall Objective Function
where $\mathcal{L}_{IM}$, $\mathcal{L}_{CM}$, $\mathcal{L}^{v}_{OCLR}$ and $\mathcal{L}^{r}_{OCLR}$ are described in detail in Sec.3.2 and Sec.3.3.
@@ 4 Experiments
4.1 Dataset and Evaluation Protocol
Datasets. Our proposed method is evaluated on two public visible-infrared datasets, namely SYSU-MM01 Wu et al (2017) and RegDB Nguyen et al (2017). SYSU-MM01 comprises 395 identities, with 22258 visible and 11909 infrared images captured by indoor and outdoor cameras. The dataset is divided into the training set (296 IDs), the validation set (99 IDs), and the test set (96 IDs). RegDB is a smaller dataset obtained from a pair of aligned visible and infrared cameras Nguyen et al (2017). It contains 412 identities, where each identity has 10 visible images and 10 infrared images.
Evaluation Metrics. All experiments follow the common evaluation protocols used for VI-ReID Ye et al (2021b, 2020, a). The evaluation metrics include Cumulative Matching Characteristic (CMC), Mean Average Precision (mAP), and Mean Inverse Negative Penalty (mINP Ye et al (2021b)). For the SYSU-MM01 dataset, following Wu et al (2021); Lu et al (2020); Mao et al (2017), we evaluate the proposed method under two search modes: the All-search mode and the Indoor-search mode. In All-search mode, the gallery set contains all visible images from both indoor and outdoor cameras, while in indoor-search mode, the gallery set only contains images from indoor cameras. Additionally, we conduct experiments under both single-shot and multi-shot settings. We repeat the evaluation 10 times with the random split of the query set and the gallery set and report the average performance. As for the RegDB dataset, we evaluate our method on two testing modes: Visible-to-Infrared and Infrared-to-Visible. The dataset is randomly split into training (206 identities) and testing (206 identities) sets. We randomly selected 206 identities for training and the remaining 206 for testing.
Implementation Details. Our proposed method is implemented using PyTorch. we adopt two-stream ResNet50 He et al (2016) pretrained on ImageNet Deng et al (2009) as our backbone. In a mini-batch, the number of classes $P$ and instances for each class $K$ are both 12. All the input images are resized to $288\times 144$. The augmentations for images are following Yang et al (2022). Besides, Linear Transform Generator (LTG Tan et al (2023)) and ColorJitter augmentations are adopted in the DCL training stage to promote the encoder to extract color-independent features. We train for a total of 90 epochs. The DCL framework is trained in the first 40 epochs, while our proposed framework is trained in the other 50 epochs. We use the Adam optimizer for training the model with weight decay 5e-4. The initial learning rate is set to 3.5e-4 and decays 10 times at the 20th, 60th, and 80th epochs. Following Cheng et al (2023a); Yang et al (2022); Wu and Ye (2023), the momentum factor $\mu$ is 0.1, and the temperature $\tau$ is 0.05, while the maximum distance for DBSCAN is set to 0.6 on SYSU-MM01 and 0.3 on RegDB. The parameter $\kappa$ for $\kappa$-reciprocal nearest neighbors in Eq.1 and Eq.2 is set to 30 following Zhong et al (2017). The hyper-parameter $\lambda$ in Eq.10 for OT is set to 25. The trade-off parameter $\alpha$ in Eq.14 and Eq.15 is set to 0.2 and $\beta$ in Eq.21 is set to 0.7.
Table 2: Comparison with the state-of-the-art methods under the Multi-shot setting on SYSU-MM01.
Method SYSU-MM01 (Multi-shot) All-search Indoor-search R1 mAP R1 mAP Supervised VI-ReID methods Zero-Pad Wu et al (2017) 19.13 10.89 24.43 18.86 cmGAN Dai et al (2018) 31.49 22.27 37.00 32.76 AlignGAN Mao et al (2017) 51.50 33.90 57.10 45.30 cm-SSFT Lu et al (2020) 63.40 62.00 73.00 72.40 MPANet Wu et al (2021) 75.58 62.91 84.22 75.11 FMCNet Zhang et al (2022a) 73.44 56.06 78.86 63.82 CIFT Li et al (2022b) 79.74 75.56 88.32 86.42 PartMix Kim et al (2023) 80.54 69.84 87.99 79.95 Unsupervised VI-ReID methods H2H Liang et al (2021) 30.31 19.12 - - DFC Si et al (2023) 44.12 28.36 - - MBCCM Cheng et al (2023a) 57.73 39.78 62.87 52.80 CHCR Pang et al (2023) 50.12 42.17 - - Ours 71.35 52.18 76.99 64.03
@@ 4.1 Dataset and Evaluation Protocol
Index Method SYSU-MM01 RegDB All-search Indoor-search Visible-to-Infrared Visible-to-Infrared R1 mAP mINP R1 mAP mINP R1 mAP mINP R1 mAP mINP 1 Baseline 41.76 38.57 24.32 45.83 54.44 50.00 50.58 48.78 35.62 51.17 47.24 33.80 The following methods are all based on the MULT module 2 Baseline + $\mathcal{L}_{CM}$ 60.02 54.92 39.18 60.66 66.92 62.24 83.74 75.55 58.26 85.29 75.08 55.50 3 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{CM}$ 60.18 55.52 39.86 60.48 67.39 62.94 85.87 76.08 58.05 84.13 74.36 55.03 4 Baseline + $\mathcal{L}_{OCLR}$ 54.34 51.03 36.21 55.85 62.68 58.12 85.87 78.72 65.41 87.86 79.48 61.70 5 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{OCLR}$ 62.35 58.22 43.10 64.77 70.89 66.38 88.40 81.20 67.08 88.93 80.27 62.73 6 Baseline + $\mathcal{L}_{CM}$ + $\mathcal{L}_{OCLR}$ 62.70 58.25 43.11 64.06 70.50 66.11 87.77 80.07 66.40 88.16 79.31 62.68 7 Baseline + $\mathcal{L}_{IM}$ + $\mathcal{L}_{CM}$ + $\mathcal{L}_{OCLR}$ 64.77 59.23 43.46 65.34 71.46 67.83 89.95 82.09 67.29 90.78 82.25 65.38
4.2 Comparison with State-of-the-Art Methods
To demonstrate the efficiency of our proposed methods, we compare our method with state-of-the-art methods under three relevant settings on SYSU-MM01 and RegDB datasets, $i.e.$, supervised VI-ReID methods, unsupervised single-modality ReID methods, and unsupervised VI-ReID methods. The results are shown in Tab.1 and Tab.2.
Comparison with Supervised VI-ReID Methods. Our proposed method achieves competitive performance with supervised method VCD+VML Tian et al (2021) on SYSU-MM01, and outperform several supervised methods including Zero-Pad Wu et al (2017), AlignGAN Mao et al (2017), cm-SSFT Mao et al (2017), DDAG Ye et al (2020) and AGW Ye et al (2021b). Our method surpasses most supervised VI-ReID methods on the RegDB dataset and demonstrates excellent performance close to state-of-the-art methods. The results reveal the significant developmental potential of unsupervised VI-ReID, yet there remains a considerable performance gap compared to the state-of-the-art supervised VI-ReID methods.
Comparison with Conventional Unsupervised Single-Modality ReID Methods. The results in Tab.1 indicate that unsupervised single-modality methods cannot effectively address the large modality-gap in cross-modality scenarios Our method outperforms the state-of-the-art unsupervised single-modality ReID method Cluster-Contrast Dai et al (2023) by a large margin of 37.23% mAP and 44.61% Rank-1 on SYSU-MM01, emphasizing the necessity of proposing cross-modality label association strategies for VI-ReID.
Comparison with Unsupervised VI-ReID Methods. As reported in Tab.1, our method outperforms the state-of-the-art GUR Yang et al (2023) by 2.24% mAP, 3.82% Rank-1 on SYSU-MM01 (All-search) under the single-shot setting, and 11.86% mAP, 16.04% Rank-1 on RegDB (Visible-to-infrared). As shown in Tab.2, our method showcases remarkable performance, surpassing the state-of-the-art CHCR Pang et al (2023) by 10.01% mAP and 21.23% Rank-1 on SYSU-MM01 under the multi-shot setting (All-search). (Note that PGM Wu and Ye (2023), GRU Yang et al (2023), among others, do not provide results for the multi-shot setting on SYSU-MM01.) The proposed method has three main advantages: (1) Unlike H2H Liang et al (2021) and OTLA Wang et al (2022), our method does not require additional datasets or camera labels for training. (2) Our MULT associates cross-modality labels from a novel structural consistency perspective and can be applied to other unsupervised cross-domain tasks. (3) Our OCLR is a plug-and-play module that can be employed in other unsupervised memory-based methods to handle label noise.
4.3 Ablation Study
To validate the effectiveness of each component of our method, we conduct ablation experiments on SYSU-MM01 and RegDB datasets, as shown in Tab.3. Our baseline consists of a two-stream ResNet trained under the DCL framework Yang et al (2022). All experiments are conducted with our MULT module. We do not perform an ablation study on $\mathcal{L}_{IM}$ as it does not address the problem of modality discrepancy.
Effectiveness of MULT. The MULT module delivers a +16.35% mAP and +18.26% Rank-1 improvement by directly incorporating $\mathcal{L}_{CM}$ (see $1^{st}$ row and $2^{nd}$ row in Tab.3) and achieves +20.66% mAP and +23.01% Rank-1 improvement when trained within our entire AMIRL framework on SYSU-MM01 (see $1^{st}$ row and $7^{th}$ row in Tab.3). To further evaluate the effectiveness of MULT, we compare MULT with other cross-modality label association methods in Tab.4, including PGM Wu and Ye (2023), MBCCM Cheng et al (2023a), CLU Yang et al (2023) and DOTLA Cheng et al (2023b). Specifically, we integrate the above methods into our AMIRL framework. The results demonstrate our MULT provides higher-quality associations than other methods, achieving performance exceeding DOTLA Cheng et al (2023b) by 4.19% mAP without OCLR, and 3.88% mAP by integrating our OCLR module. The core effectiveness of our MULT lies in the generated soft pseudo-labels containing richer contextual structural information within the feature space, which promotes the network’s learning of shared fine-grained features both within clusters and across clusters.